OpenStudy (anonymous):

What is the area of a circle with a diameter of twenty-eight meters? Use twenty-two sevenths for pi.

OpenStudy (johnweldon1993):

Well you know the area of a circle is \[\pi r^2\] We have the diameter.....but we need the radius for the equation...so divide 28 in half to get 14...now have have all we need since we know to use 22/7 for pi \[Area = \frac{ 22 }{ 7 } \times 14^2\]

OpenStudy (anonymous):



OpenStudy (johnweldon1993):

Lol